OpenStudy (anonymous):

solve the equation 5(10-x)-2(3x+5=-37

OpenStudy (johnweldon1993):

5(10 - x) - 2(3x + 5) = -37 we need to distribute first...do you know how to do this?

OpenStudy (anonymous):

yes... I got all the way but the answer I got doesn't make sense on the assignment I have to complete.

OpenStudy (anonymous):

I got -3/11

OpenStudy (anonymous):

x=-3/11

OpenStudy (johnweldon1993):

Well here let's see what you did...mind posting the steps you took?

OpenStudy (anonymous):

distributed the 5 first to equal... 50-5x distributed the -2 to equal.... -6x-10 I combined like terms... 40-11x=-37 And now by writing it all down I understand what I did wrong... I forgot about the negative sign on the 37 that's why the answer wasn't making sense... so now its ...x=7

OpenStudy (johnweldon1993):

That is correct! good job for catching your mistake!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks!!!

OpenStudy (johnweldon1993):

no problem!